# for functionals in the $\mathcal{A}$-free setting

### v1

@@ 1. Introduction
When $f$ is a Carathéodory function satisfying the usual $p$-growth conditions (see (2.6)), the study of the minimization of the integral functional
@@ 2. Notation and preliminaries
Throughout this paper, $N,d,l\in\mathbb{N}$ and $c_{0},c_{1},p,q\in(1,+\infty)$ are fixed, with $\frac{1}{p}+\frac{1}{q}=1$ and $c_{1}\geq c_{0}\max\{2^{p-2},1\}$. We define $\mathcal{O}(\mathbb{R}^{N})$ to be the collection of all bounded open subsets of $\mathbb{R}^{N}$ and, for $D\in\mathcal{O}(\mathbb{R}^{N})$, we define $\mathcal{O}(D):=\{D\cap U\colon U\in\mathcal{O}(\mathbb{R}^{N})\}$ to be the collection of all open subsets of $D$. For every $x\in\mathbb{R}^{N}$ and $\rho>0$, we consider the open cube $Q_{\rho}(x)$ with center $x$, side length $\rho$, and sides parallel to the coordinate axes.
@@
Next, we introduce the class of integrands that we consider in our analysis.
@@
Let $\mathcal{F}$ be the collection of all Carathéodory functions $f\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,+\infty)$ satisfying, for every $x\in\mathbb{R}^{N}$ and every $\xi,\xi_{1},\xi_{2}\in\mathbb{R}^{d}$, the following $p$-growth and $p$-Lipschitz continuity conditions:
$\frac{1}{c_{0}}|\xi|^{p}-c_{0}\leq f(x,\xi)\leq c_{0}(1+|\xi|^{p}),$ (2.6)
@@
where $a\land b:=\min\{a,b\}$. Finally, let $\mathcal{F}_{\mathrm{qc}}$ be the collection of all functions $f\in\mathcal{F}$ such that $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for a.e. $x\in\mathbb{R}^{N}$.
@@
Remark 2.4.
There exists a constant $c_{2}$, depending only on $c_{0}$, $c_{1}$, and $p$, such that (2.6) and (2.7) imply
$|f(x,\xi_{1})-f(x,\xi_{2})|\leq c_{2}(1+|\xi_{1}|^{p-1}+|\xi_{2}|^{p-1})|\xi_{1}-\xi_{2}|$ (2.8)
for every $x\in\mathbb{R}^{N}$ and every $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$. Conversely, if (2.6) and (2.8) hold, then there exists a constant $c_{3}$, depending only on $c_{2}$ and $p$, such that
$|f(x,\xi_{1})-f(x,\xi_{2})|\leq c_{3}\big{(}1+(|\xi_{1}|\land|\xi_{2}|)^{p-1}+|\xi_{1}-\xi_{2}|^{p-1}\big{)}|\xi_{1}-\xi_{2}|,$
which implies that there exists a constant $c_{1}$, depending only on $c_{0}$, $c_{3}$, and $p$, such that (2.7) holds. Our preference for (2.7) is due to the fact that, unlike (2.8), it is stable under $\Gamma$-convergence.
@@
Let $\{L^{p},\mathcal{O}\}$ be the set of pairs $(u,D)$ with $u\in L^{p}(D,\mathbb{R}^{d})$ and $D\in\mathcal{O}(\mathbb{R}^{N})$, and let $\mathcal{I}$ be the collection of all functionals $F\colon\{L^{p},\mathcal{O}\}\to[0,+\infty)$ satisfying the following properties for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u,v\in L^{p}(D;\mathbb{R}^{d})$:
$\thinspace(a)\,\,\tfrac{1}{c_{0}}\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p}-c_{0}|D|\leq F(u,D)\leq c_{0}\big{(}|D|+\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p}\big{)},$ (2.10)
$\thinspace(b)\,\,|F(u,D)-F(v,D)|\leq c_{1}\big{(}|D|^{\frac{p-1}{p}}+(F(u,D)\land F(v,D))^{\frac{p-1}{p}}+\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}\big{)}\|u-v\|_{L^{p}(D;\mathbb{R}^{d})},\quad$ (2.11)
[…]
Here, and henceforth, if $u\in L^{p}(D;\mathbb{R}^{d})$ and $B\in\mathcal{O}(D)$, we simply write $F(u,B)$ instead of $F(u|_{B},B)$. Finally, let $\mathcal{I}_{\mathrm{sc}}$ be the collection of all functionals $F\in\mathcal{I}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the functional $u\mapsto F(u,D)$ is lower semicontinuous for the topology induced on $L^{p}(D;\mathbb{R}^{d})$ by the norm $\|\cdot\|^{\mathcal{A}}_{D}$.
@@
Let $F\in\mathcal{I}$. By the lower bound in (2.10), if $D\in\mathcal{O}(\mathbb{R}^{N})$ and $(u_{k})_{k\in\mathbb{N}}$ is a sequence in $L^{p}(D;\mathbb{R}^{d})$ converging to $u\in L^{p}(D;\mathbb{R}^{d})$ in the topology induced by the norm $\|\cdot\|^{\mathcal{A}}_{D}$ and such that $(F(u_{k},D))_{k\in\mathbb{N}}$ is bounded, then $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$. Therefore, $F\in\mathcal{I}_{\mathrm{sc}}$ if and only if we have for every $D\in\mathcal{O}(\mathbb{R}^{N})$ that
@@
for every $u\in L^{p}(D;\mathbb{R}^{d})$ and every sequence $(u_{k})_{k\in\mathbb{N}}\subset(L^{p}(D;\mathbb{R}^{d})$ such that $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$ and $\mathcal{A}_{D}u_{k}\to\mathcal{A}u$ strongly in $W^{-1,p}(D;\mathbb{R}^{l})$.
@@
belongs to $\mathcal{I}$. Indeed, (2.11) follows from (2.7) by using Hölder’s inequality. If $f\in\mathcal{F}_{\mathrm{qc}}$, then $F\in\mathcal{I}_{\mathrm{sc}}$ by [19, Theorem 3.7] and Remark 2.7.
@@
By analogy with Remark 2.4, there exists a constant $c_{2}$, depending only on $c_{0}$, $c_{1}$, and $p$, such that if $F$ satisfies (2.10) and (2.11), then
$|F(u,D)-F(v,D)|\leq c_{2}\big{(}|D|^{\frac{p-1}{p}}+\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}+\|v\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}\big{)}\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}$ (2.14)
for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u,v\in L^{p}(D;\mathbb{R}^{d})$. Conversely, if (2.10) and (2.14) hold, then there exists a constant $c_{3}$, depending only on $c_{0}$, $c_{2}$, and $p$, such that
$|F(u,D)-F(v,D)|\leq c_{3}\big{(}|D|^{\frac{p-1}{p}}+(F(u,D)\land F(v,D))^{\frac{p-1}{p}}+\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}\big{)}\|u-v\|_{L^{p}(D;\mathbb{R}^{d})},$
which implies that (2.11) holds with $c_{1}$ replaced by $c_{3}$. As before, our preference for (2.11) is due to the fact that, unlike (2.14), it is stable under $\Gamma$-convergence.
@@
Similarly to Remark 2.5, we see by exchanging the roles of $u$ and $v$ that (2.11) is equivalent to
@@ 3. $\Gamma$-convergence in the unconstrained setting
Since $\nu$ is a finite measure, we can select $\eta\in(0,\tfrac{\tau}{2})$ such that and $\nu(\partial D^{\eta})=0$ (see [17, Proposition 1.15]). For every $m\in\mathbb{N}$ with $0<\frac{1}{m}<\frac{\tau}{2}$, we define the sets
@@
Indeed, the definition of $F^{\prime\prime}(\cdot,D)$ and (2.10) yield
@@
which exist by [11, Proposition 8.1]. By Lemma 3.2, there exists a sequence $(w_{k})_{k\in\mathbb{N}}\subset L^{p}(D_{1}\cup B;\mathbb{R}^{d})$ such that (3.3) and (3.4) hold, and so
@@
Step 4 (Compactness property). Let $\mathcal{D}$ be the countable collection of the open sets that are finite unions of open rectangles with rational vertices. Using the compactness of $\Gamma$-convergence on separable metric spaces (see [11, Theorem 8.5]) and a diagonal argument, we obtain a subsequence of $(F_{k})_{k\in\mathbb{N}}$, which we do not relabel, for which
@@
Step 6 (Proof of (2.10) for $F$). Fix $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u\in L^{p}(D;\mathbb{R}^{d})$. By Step 5, there exists a sequence $(u_{k})_{k\in\mathbb{N}}\subset L^{p}(D;\mathbb{R}^{d})$ such that
@@
Then, using the lower bound in (2.10) for $F_{k}$, we have for all sufficiently large $k\in\mathbb{N}$ that
@@
The preceding estimate and (3.14) yield that $(u_{k})_{k\in\mathbb{N}}$ is a bounded sequence in $L^{p}(D;\mathbb{R}^{d})$. Hence, $u_{k}\to u$ weakly in $L^{p}(D;\mathbb{R}^{d})$, and so
@@
Finally, (3.18) and (3.14) show that (2.10) holds for $F$.
Step 7 (Proof of (2.11) for $F$). Fix $D\in\mathcal{O}(\mathbb{R}^{N})$. By (2.15) for $F_{k}$, we have
@@
This inequality, together with (3.20) and (3.21), leads to (2.15) for $F$. As observed in Remark 2.10, we then conclude that (2.11) holds for $F$.
Step 8 (Proof of (2.12) for $F$). Fix $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u\in L^{p}(D;\mathbb{R}^{d})$, and let $\alpha\colon\mathcal{O}(D)\to[0,+\infty)$ be the (increasing) set function defined by setting $\alpha(B):=F(u,B)$ for every $B\in\mathcal{O}(D)$. Invoking [11, Theorem 14.23], property (2.12) is satisfied provided that $\alpha$ is subadditive, superadditive, and inner regular in $\mathcal{O}(D).$ The inner regularity holds by (3.17), while the simple proof of the superadditivity can be obtained as in [11, Proposition 16.12]. We are then left to show that $\alpha$ is superadditive in $\mathcal{O}(D)$, which amounts to proving that
$\alpha(B_{1}\cup B_{2})-\delta<F^{\prime\prime}(u,B^{\prime}).$
Let $B^{\prime}_{1},B^{\prime\prime}_{1},B_{2}^{\prime}\in\mathcal{O}(D)$ be such that $B_{1}^{\prime}\subset\subset B^{\prime\prime}_{1}\subset\subset B_{1}$, $B_{2}^{\prime}\subset\subset B_{2}$, and $B^{\prime}\subset\subset B_{1}^{\prime}\cup B^{\prime}_{2}$ (see [11, Lemma 14.20] for instance). Then, using (3.9), (3.17), and the monotonicity of $F^{\prime\prime}(u,\cdot)$ proved in Step 1, we obtain
from which (3.22) follows by letting $\delta\to 0$. ∎
We now prove that every functional in $\mathcal{I}$ can be represented by an integral whose integrand belongs to $\mathcal{F}$.
@@
Let $F\in\mathcal{I}$. For every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$, we set
@@
for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u\in L^{p}(D,\mathbb{R}^{d})$. If $F\in\mathcal{I}_{\mathrm{sc}}$, then the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for every $x\in\mathbb{R}^{N}$; that is, $f\in\mathcal{F}_{\mathrm{qc}}$.
@@
By (2.10) and (3.23), the function $f$ satisfies (2.6). Since $F$ verifies (2.15), we deduce that(2.9) holds for $f$, which is equivalent to (2.7) (see Remark 2.4).
Let us fix $D\in\mathcal{O}(\mathbb{R}^{N})$. By (2.12), the set function $B\mapsto F(\xi,B)$ defined on $\mathcal{O}(D)$ can be extended to a measure defined on all Borel subsets of $D$. By (2.10), this measure is absolutely continuous with respect to the Lebesgue measure. By (3.23) and by the Lebesgue Differentiation Theorem, the function $x\mapsto f(x,\xi)$ is measurable on $\mathbb{R}^{N}$ for every $\xi\in\mathbb{R}^{d}$, and

### v2

@@ 1. Introduction
When $f$ is a Carathéodory function satisfying the usual $p$-growth condition (see (2.10)), the study of the minimization of the integral functional
@@ 2. Notation and preliminaries
Throughout this paper, $N,d,l\in\mathbb{N}$ and $p,q\in(1,+\infty)$ are fixed, with $\frac{1}{p}+\frac{1}{q}=1$. We define $\mathcal{O}(\mathbb{R}^{N})$ to be the collection of all bounded open subsets of $\mathbb{R}^{N}$ and, for $D\in\mathcal{O}(\mathbb{R}^{N})$, we define $\mathcal{O}(D):=\{D\cap U\colon U\in\mathcal{O}(\mathbb{R}^{N})\}$ to be the collection of all open subsets of $D$. For every $x\in\mathbb{R}^{N}$ and $\rho>0$, we consider the open cube $Q_{\rho}(x)$ with center $x$, side length $\rho$, and sides parallel to the coordinate axes.
@@
The wave cone (or characteristic cone) $\Lambda$ associated with the operator $\mathcal{A}$ is the subset of $\mathbb{R}^{d}$ defined by
$\Lambda:=\bigcup_{w\in\mathbb{R}^{N}\setminus\{0\}}\ker\left(\sum_{i=1}^{N}A^{i}w_{i}\right).$ (2.6)
@@
Remark 2.3.
If $g$ is $\mathcal{A}$-quasiconvex and satisfies the $p$-growth condition
$\frac{1}{c_{0}}|\xi|^{p}-c_{0}\leq g(\xi)\leq c_{0}(1+|\xi|^{p})\quad\hbox{for every }\xi\in\mathbb{R}^{d}$ (2.7)
for some constant $c_{0}\geq 1$, then $g$ is $\Lambda$-convex on $\mathbb{R}^{d}$, i.e.,
$g(\theta\xi_{1}+(1-\theta)\xi_{2})\leq\theta g(\xi_{1})+(1-\theta)g(\xi_{2})$
for every $\xi_{1}$, $\xi_{2}\in\mathbb{R}^{d}$ with $\xi_{1}-\xi_{2}\in\Lambda$ and every $\theta\in[0,1]$ (see, e.g., the proof of [4, Lemma 2.19]). If the vector space $\mathrm{span}(\Lambda)$ generated by the wave cone $\Lambda$ coincides with $\mathbb{R}^{d}$, then $\Lambda$-convexity and (2.7) imply that there exists a constant $\underline{c}_{1}>0$, depending only on $c_{1}$, $p$, and $d$, such that the $p$-Lipschitz condition
$|g(\xi_{1})-g(\xi_{2})|\leq\underline{c}_{1}(1+|\xi_{1}|^{p-1}+|\xi_{2}|^{p-1})|\xi_{1}-\xi_{2}|$ (2.8)
holds for every $\xi_{1}$, $\xi_{2}\in\mathbb{R}^{d}$ (see, e.g., [23, Lemma 2.3]).
The equality $\mathrm{span}(\Lambda)=\mathbb{R}^{d}$ is satisfied in many interesting cases, for instance when $\mathcal{A}=\operatorname{curl}$ or $\mathcal{A}=\operatorname{div}$ (see [20, Remark 3.3] for the precise definitions).
Remark 2.4.
For functions $g\colon\mathbb{R}^{d}\to[0,+\infty)$, the $p$-Lipschitz condition can be expressed in many equivalent ways. Indeed, if $g$ satisfies (2.8), then we can find a constant $\bar{c}_{1}>0$, depending only on $\underline{c}_{1}$ and $p$, such that
$|g(\xi_{1})-g(\xi_{2})|\leq\bar{c}_{1}\big(1+(|\xi_{1}|\land|\xi_{2}|)^{p-1}+|\xi_{1}-\xi_{2}|^{p-1}\big)|\xi_{1}-\xi_{2}|$
holds for every $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$, where $a\land b:=\min\{a,b\}$. If $g$ satisfies also (2.7), the previous inequality implies that there exists a constant $\hat{c}_{1}\geq 1$, depending only on $c_{0}$, $\bar{c}_{1}$, and $p$, such that
$|g(\xi_{1})-g(\xi_{2})|\leq\hat{c}_{1}\big(1+(g(\xi_{1})\land g(\xi_{2}))^{\frac{p-1}{p}}+|\xi_{1}-\xi_{2}|^{p-1}\big)|\xi_{1}-\xi_{2}|$ (2.9)
holds for every $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$.
Conversely, if $g$ satisfies (2.7) and (2.9), we can find a constant $\check{c}_{1}>0$, depending only on $c_{0}$, $\hat{c}_{1}$, and $p$, such that (2.8) holds with $\underline{c}_{1}$ replaced by $\check{c}_{1}$.
In the rest of the paper, we prefer to express the $p$-Lipschitz condition in the rather unusual form (2.9), because in this inequality the constant is stable under $\Gamma$-convergence, while this is not true for the more usual form (2.8).
Next, we introduce the class of integrands that we consider in our analysis. Throughout the paper $c_{0}$ and $c_{1}$ are fixed constants with $c_{0}\geq 1$ and $c_{1}\geq\hat{c}_{1}\geq 1$, where $\hat{c}_{1}$ is the constant in (2.9) corresponding to the constant $\bar{c}_{1}$ obtained from $\underline{c}_{1}$ in (2.8).
@@
Let $\mathcal{F}$ be the collection of all Carathéodory functions $f\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to[0,+\infty)$ satisfying the following $p$-growth condition:
$\frac{1}{c_{0}}|\xi|^{p}-c_{0}\leq f(x,\xi)\leq c_{0}(1+|\xi|^{p})$ (2.10)
for every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$. Let $\mathcal{F}_{\mathrm{Lip}}$ be the collection of all $f\in\mathcal{F}$ such that
@@
for every $x\in\mathbb{R}^{N}$ and $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$. Finally, let $\mathcal{F}_{\mathrm{qc}}$ be the collection of all functions $f\in\mathcal{F}$ such that $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for every $x\in\mathbb{R}^{N}$.
We observe that there is no loss of generality in the constraint on the constant $c_{1}$ in the preceding definition because if (2.11) holds for a constant $c_{1}$, then it also holds for any constant $c\geq c_{1}$.
@@
Remark 2.7.
As discussed in Remark 2.4, there exists a constant $\check{c}_{1}>0$, depending only on $c_{0}$, $c_{1}$, and $p$, such that (2.10) and (2.11) imply
$|f(x,\xi_{1})-f(x,\xi_{2})|\leq\check{c}_{1}(1+|\xi_{1}|^{p-1}+|\xi_{2}|^{p-1})|\xi_{1}-\xi_{2}|$ (2.13)
for every $x\in\mathbb{R}^{N}$ and every $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$.
Remark 2.8.
By Remarks 2.3 and 2.4, in the case where the vector space $\mathrm{span}(\Lambda)$ generated by the wave cone $\Lambda$ coincides with $\mathbb{R}^{d}$, we have that every $f\in\mathcal{F}_{\mathrm{qc}}$ satisfies (2.11) with $\hat{c}_{1}$ in place of $c_{1}$. Since we are assuming throughout that $\hat{c}_{1}\leq c_{1}$, we conclude that $\mathcal{F}_{\mathrm{qc}}\subset\mathcal{F}_{\mathrm{Lip}}$ in this case.
@@
Let $\{L^{p},\mathcal{O}\}$ be the set of pairs $(u,D)$ with $u\in L^{p}(D,\mathbb{R}^{d})$ and $D\in\mathcal{O}(\mathbb{R}^{N})$, and let $\mathcal{I}$ be the collection of all functionals $F\colon\{L^{p},\mathcal{O}\}\to[0,+\infty)$ satisfying the following properties for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u\in L^{p}(D;\mathbb{R}^{d})$:
  • (a)
    $\tfrac{1}{c_{0}}\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p}-c_{0}|D|\leq F(u,D)\leq c_{0}\big(|D|+\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p}\big),$
  • (b)
    the set function $B\mapsto F(u,B)$ defined for every $B\in\mathcal{O}(D)$ can be extended to a nonnegative measure defined on all Borel subsets of $D$.
Here, and henceforth, if $u\in L^{p}(D;\mathbb{R}^{d})$ and $B\in\mathcal{O}(D)$, we simply write $F(u,B)$ instead of $F(u|_{B},B)$.
Let $\mathcal{I}_{\mathrm{Lip}}$ be the collection of all functionals $F\in\mathcal{I}$ such that
  • (c)
    $|F(u,D)-F(v,D)|\leq c_{1}\big(|D|^{\frac{p-1}{p}}+(F(u,D)\land F(v,D))^{\frac{p-1}{p}}+\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}\big)\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}$
for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u$, $v\in L^{p}(D;\mathbb{R}^{d})$.
Finally, let $\mathcal{I}_{\mathrm{sc}}$ be the collection of all functionals $F\in\mathcal{I}$ such that for every $D\in\mathcal{O}(\mathbb{R}^{N})$, the functional $u\mapsto F(u,D)$ is lower semicontinuous for the topology induced on $L^{p}(D;\mathbb{R}^{d})$ by the norm $\|\cdot\|^{\mathcal{A}}_{D}$.
@@
Let $F\in\mathcal{I}$. By the lower bound in (a) of Definition 2.9, if $D\in\mathcal{O}(\mathbb{R}^{N})$ and $(u_{k})_{k\in\mathbb{N}}$ is a sequence in $L^{p}(D;\mathbb{R}^{d})$ converging to $u\in L^{p}(D;\mathbb{R}^{d})$ in the topology induced by the norm $\|\cdot\|^{\mathcal{A}}_{D}$ and such that $(F(u_{k},D))_{k\in\mathbb{N}}$ is bounded, then $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$. Therefore, $F\in\mathcal{I}_{\mathrm{sc}}$ if and only if we have for every $D\in\mathcal{O}(\mathbb{R}^{N})$ that
@@
for every $u\in L^{p}(D;\mathbb{R}^{d})$ and every sequence $(u_{k})_{k\in\mathbb{N}}\subset L^{p}(D;\mathbb{R}^{d})$ such that $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$ and $\mathcal{A}_{D}u_{k}\to\mathcal{A}u$ strongly in $W^{-1,p}(D;\mathbb{R}^{l})$. In particular, this inequality holds whenever $u_{k}\to u$ strongly in $L^{p}(D;\mathbb{R}^{d})$. Hence, for every $F\in\mathcal{I}_{\mathrm{sc}}$ and $D\in\mathcal{O}(\mathbb{R}^{N})$ the functional $F(\cdot,D)$ is lower semicontinuous in the strong topology of $L^{p}(D;\mathbb{R}^{d})$.
@@
belongs to $\mathcal{I}$. If $f\in\mathcal{F}_{\mathrm{Lip}}$, then $F\in\mathcal{I}_{\mathrm{Lip}}$. Indeed, condition (c) of Definition 2.9 follows from (2.11) by using Hölder’s inequality. If $f\in\mathcal{F}_{\mathrm{qc}}$, then $F\in\mathcal{I}_{\mathrm{sc}}$ by [20, Theorem 3.7] and Remark 2.10.
@@
By analogy with Remark 2.4, there exists a constant $\check{c}_{1}$, depending only on $c_{0}$, $c_{1}$, and $p$, such that if $F\in\mathcal{I}_{\mathrm{Lip}}$, then
$|F(u,D)-F(v,D)|\leq\check{c}_{1}\big(|D|^{\frac{p-1}{p}}+\|u\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}+\|v\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}\big)\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}$ (2.15)
for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u,v\in L^{p}(D;\mathbb{R}^{d})$. Conversely, if the growth conditions (a) of Definition 2.9 and the Lipschitz condition (2.15) hold, then there exists a constant $\hat{c}_{1}$, depending only on $c_{0}$, $\check{c}_{1}$, and $p$, such that
$|F(u,D)-F(v,D)|\leq\hat{c}_{1}\big(|D|^{\frac{p-1}{p}}+(F(u,D)\land F(v,D))^{\frac{p-1}{p}}+\|u-v\|_{L^{p}(D;\mathbb{R}^{d})}^{p-1}\big)\|u-v\|_{L^{p}(D;\mathbb{R}^{d})},$
which implies that the Lipschitz condition (c) of Definition 2.9 holds with $c_{1}$ replaced by $\hat{c}_{1}$. As before, our preference for condition (c) of Definition 2.9 is due to the fact that, unlike (2.15), it is stable under $\Gamma$-convergence.
@@
Similarly to Remark 2.6, we see by exchanging the roles of $u$ and $v$ that the Lipschitz condition (c) of Definition 2.9 is equivalent to the inequality
@@ 3. $\Gamma$-convergence in the unconstrained setting
If, in addition, $F_{k}\in\mathcal{I}_{\mathrm{Lip}}$ for every $k\in\mathbb{N}$, then $F\in\mathcal{I}_{\mathrm{Lip}}$.
@@
Since $\nu$ is a finite measure, we can select $\eta\in(0,\tfrac{\tau}{2})$ such that and $\nu(\partial D^{\eta})=0$ (see [18, Proposition 1.15]). For every $m\in\mathbb{N}$ with $0<\frac{1}{m}<\frac{\tau}{2}$, we define the sets
@@
Indeed, the definition of $F^{\prime\prime}(\cdot,D)$ and the upper bound in (a) of Definition 2.9 yield
@@
which exist by [12, Proposition 8.1]. By Lemma 3.2, there exists a sequence $(w_{k})_{k\in\mathbb{N}}\subset L^{p}(D_{1}\cup B;\mathbb{R}^{d})$ such that (3.3) and (3.4) hold, and so
@@
Step 4 (Compactness property). Let $\mathcal{D}$ be the countable collection of the open sets that are finite unions of open rectangles with rational vertices. Using the compactness of $\Gamma$-convergence on separable metric spaces (see [12, Theorem 8.5]) and a diagonal argument, we obtain a subsequence of $(F_{k})_{k\in\mathbb{N}}$, which we do not relabel, for which
@@
Step 6 (Proof of condition (a) of Definition 2.9 for $F$). Fix $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u\in L^{p}(D;\mathbb{R}^{d})$. By Step 5, there exists a sequence $(u_{k})_{k\in\mathbb{N}}\subset L^{p}(D;\mathbb{R}^{d})$ such that
@@
Then, using the lower bound in condition (a) of Definition 2.9 for $F_{k}$, we have for all sufficiently large $k\in\mathbb{N}$ that
@@
The preceding estimate and (3.14) yield that $(u_{k})_{k\in\mathbb{N}}$ is a bounded sequence in $L^{p}(D;\mathbb{R}^{d})$. Hence, $u_{k}\rightharpoonup u$ weakly in $L^{p}(D;\mathbb{R}^{d})$, and so
@@
Finally, (3.14) and (3.18) show that condition (a) of Definition 2.9 holds for $F$.
Step 7 (Proof of condition (b) of Definition 2.9 for $F$). Fix $D\in\mathcal{O}(\mathbb{R}^{N})$ and $u\in L^{p}(D;\mathbb{R}^{d})$, and let $\alpha\colon\mathcal{O}(D)\to[0,+\infty)$ be the (increasing) set function defined by setting $\alpha(B):=F(u,B)$ for every $B\in\mathcal{O}(D)$. Invoking [12, Theorem 14.23], condition (b) of Definition 2.9 is satisfied for $F$ provided that $\alpha$ is subadditive, superadditive, and inner regular in $\mathcal{O}(D).$ The inner regularity holds by (3.17), while the simple proof of the superadditivity can be obtained as in [12, Proposition 16.12]. We are then left to show that $\alpha$ is superadditive in $\mathcal{O}(D)$, which amounts to proving that
$\alpha(B_{1}\cup B_{2})-\delta<F^{\prime\prime}(u,B^{\prime}).$
Let $B^{\prime}_{1},B^{\prime\prime}_{1},B_{2}^{\prime}\in\mathcal{O}(D)$ be such that $B_{1}^{\prime}\subset\subset B^{\prime\prime}_{1}\subset\subset B_{1}$, $B_{2}^{\prime}\subset\subset B_{2}$, and $B^{\prime}\subset\subset B_{1}^{\prime}\cup B^{\prime}_{2}$ (see [12, Lemma 14.20] for instance). Then, using (3.9), (3.17), and the monotonicity of $F^{\prime\prime}(u,\cdot)$ proved in Step 1, we obtain
from which (3.19) follows by letting $\delta\to 0$.
Step 8 (Proof of condition (c) of Definition 2.9 for $F$). Assume that $F_{k}\in\mathcal{I}_{Lip}$ for every $k\in\mathbb{N}$ and fix $D\in\mathcal{O}(\mathbb{R}^{N})$. By (2.16) for $F_{k}$, we have
@@
This inequality, together with (3.21) and (3.22), leads to (2.16) for $F$. As observed in Remark 2.13, we then conclude that condition (c) of Definition 2.9 holds for $F$. ∎
We now prove that every functional in $\mathcal{I}_{\mathrm{Lip}}$ can be represented by an integral whose integrand belongs to $\mathcal{F}_{\mathrm{Lip}}$.
@@
Let $F\in\mathcal{I}_{\mathrm{Lip}}$. For every $x\in\mathbb{R}^{N}$ and $\xi\in\mathbb{R}^{d}$, we set
@@
for every $D\in\mathcal{O}(\mathbb{R}^{N})$ and every $u\in L^{p}(D,\mathbb{R}^{d})$. If, in addition, $F\in\mathcal{I}_{\mathrm{Lip}}\cap\mathcal{I}_{\mathrm{sc}}$, then the function $\xi\mapsto f(x,\xi)$ is $\mathcal{A}$-quasiconvex for every $x\in\mathbb{R}^{N}$; that is, $f\in\mathcal{F}_{\mathrm{Lip}}\cap\mathcal{F}_{\mathrm{qc}}$.
@@
By condition (a) of Definition 2.9 and (3.23), the function $f$ satisfies (2.10). Since $F$ satisfies (2.16) (see Remark 2.13), we deduce that (2.12) holds for $f$, which is equivalent to (2.11) (see Remark 2.6).
Let us fix $D\in\mathcal{O}(\mathbb{R}^{N})$. By condition (b) of Definition 2.9 the set function $B\mapsto F(\xi,B)$ defined on $\mathcal{O}(D)$ can be extended to a measure defined on all Borel subsets of $D$. By condition (a) of Definition 2.9, this measure is absolutely continuous with respect to the Lebesgue measure. By (3.23) and by the Lebesgue Differentiation Theorem, the function $x\mapsto f(x,\xi)$ is measurable on $\mathbb{R}^{N}$ for every $\xi\in\mathbb{R}^{d}$, and